The mother is 6 years more than three times older than her son four years ago, she was 5 times older as her son. How old is her son now?
Answer by Cromlix(4381)

The mother is 6 years more than three times older than her son
M = 6 + 3S
four years ago, she was 5 times older as her son.
M - 4 = 5(S - 4)
Substitute M with 6 + 3S from equation above.
6 + 3S - 4 = 5(S - 4)
6 + 3S - 4 = 5S - 20
Collect like terms.
2S = 22
S = 11
Her son's current age is 11 years old.
Hope this helps.
:-)